OpenStudy (anonymous):

solve for x. ^4sqrtx-7 =2

OpenStudy (anonymous):



OpenStudy (anonymous):

x - 7 = 2^4 x -7 = 16 x = 23

OpenStudy (anonymous):

Taking 4th power both the sides: